Find values of x for which   x^3+1=2*(cube root of (2x-1))

Grade:12th pass

1 Answers

Vikas TU
7 years ago
  x^3+1=2*(cube root of (2x-1))
cubing both sides
 
(x^3+1)^3  = 8 (2x – 1)
x^6 + 1 + 3x^6 + 3x^3 = 8(2x – 1)
4x^6 + 3x^3  -16x + 9 = 0
 
Now make the graph and plot the graph by using wavy curve method,
U will get the approximate roots.